# A decoder-only foundation model for time-series forecasting

### v2

@@ Abstract
Author names are listed in alphabetical order.1. Google Research, Correspond to senrajat@google.com
Motivated by recent advances in large language models for Natural Language Processing (NLP), we design a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset. Our model is based on pretraining a patched-decoder style attention model on a large time-series corpus comprising both real-world and synthetic datasets. Experiments on a diverse set of previously unseen forecasting datasets suggests that the model can yield accurate zero-shot forecasts across different domains, forecasting horizons and temporal granularities.
Machine Learning, ICML
[…]
@@ 1 Introduction
Time-series data is ubiquitous in various domains such as retail, finance, manufacturing, healthcare and natural sciences. In many of these domains, one of the most important use-cases of time-series data is forecasting. Time series forecasting is critical to several scientific and industrial applications, like retail supply chain optimization, energy and traffic prediction, and weather forecasting. In recent times, Deep learning models [25, 21] have emerged as a popular approach for forecasting rich, multivariate, time series data, often outperforming classical statistical approaches such as ARIMA or GARCH [3]. In several forecasting competitions such as the M5 competition [17] and IARAI Traffic4cast contest [13], almost all the winning solutions are based on deep neural networks.
@@
This motivates the question: “Can large pretrained models trained on massive amounts of time-series data learn temporal patterns that can be useful for time-series forecasting on previously unseen datasets?” In particular, can we design a time series foundation model that obtains good zero-shot out-of-the-box forecasting performance on previously-unseen datasets? Such a time series foundation model, if possible, would bring significant benefits for downstream forecasting users in terms of significantly reduced training data and compute requirements. It is not immediately obvious that such a foundation model for time series forecasting is possible. Unlike in NLP, there is no well defined vocabulary or grammar for time-series. Additionally, the model would need to support forecasting with varying history lengths (context) , prediction lengths (horizon) and time granularities. Furthermore, unlike the huge volume of public text data for pretraining language models, vast amounts of time series data is not readily available. In spite of these issues, we provide evidence to answer the above question in the affirmative.
In particular, we design a single foundation model for time series forecasting that, when applied to a variety of previously-unseen forecasting datasets across different domains, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets). Our model can work well across different forecasting history lengths, prediction lengths and time granularities at inference time. The key elements of our foundation model are twofold: 1) a large-scale time series corpus built using both synthetic and real-world (mostly using time series data from web search queries [^1] https://trends.google.com and Wikipedia page visits [^2] https://wikimedia.org/api/rest_v1/), that meets the volume and diversity of data needed for training our foundation model, and 2) a patched-decoder style attention architecture that can be efficiently pre-trained on this time series corpus.
Compared to the latest large language models, our time series foundation model is much smaller in both parameter size (200M parameters) and pretraining data size (100B timepoints); yet we show that even at such scales it is possible to build a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time series data. Our work also suggests that unlike recent work [9] that recommends Large Language Models such as GPT-3 and LLama-2 as out-of-the-box zero-shot forecasters, foundation models trained from scratch exclusively on time-series data can obtain much better zero-shot performance at a fraction of its costs.
@@ 2 Related Work
In the last decade, deep learning models [25, 21] have emerged as powerful contenders in forecasting time-series in the presence of large training datasets and have been shown to outperform traditional statistical methods such as ARIMA, Exponential smoothing [18]. Forecasting models can be categorized broadly into: (i) Local univariate models that include traditional methods like ARIMA, exponential smoothing [18] and non-autoregressive models like Prophet [27]. These models are trained individually for each time-series in a dataset in order to predict the corresponding time-series’s future. (ii) Global univariate models like DeepAR [25], Temporal Convolutions [2], N-BEATS [21] and long-term forecasting models such as [19, 6] that are trained globally on many time-series but during inference they predict the future of a time-series as a function of its own past and other related covariates. (iii) Global multivariate models that take in the past of all time-series in the dataset to predict the future of all the time-series. Such models include the classical VAR model [37] as well as deep learning models like [26, 35, 34] to name a few.
All the works cited above have primarily been applied in the supervised setting with the notable exception of PatchTST [19] and N-BEATS [21]. PatchTST has a section on dataset to dataset transfer learning in the semi-supervised setting. The patching in our decoder-only model is inspired by [19]. [22] also show that the N-BEATS architecture lends itself to transfer learn between various source-target dataset pairs. However, none of these works aim to train a single foundation model that can work on a plethora of datasets. For a more in-depth discussion about transfer learning in time-series we refer the reader to the survey in [16].
There has been some recent work on re-using or fine-tuning large language models for time-series forecasting. In particular,  [9] benchmarks pretrained LLMs like GPT-3 and LLaMA-2 for zero-shot forecasting performance. As we show later, our model obtains much superior zero-shot performance at a tiny fraction of these model sizes. Zhou et al. [36] and [5] show how to fine-tune a GPT-2 [24] backbone model for time-series forecasting tasks. Both the works have a section on zero-shot forecasting on a target dataset after having trained on a source dataset. With the exception of a transfer-learning study (forecasting on a target dataset after having trained on a source dataset), these papers mostly focus on fine-tuning a pretrained model on a target dataset, and not on pretraining a single foundation model that shows out-of-the box zero-shot performance on a variety of datasets. To the best of our knowledge, the very recent work in TimeGPT-1 [7] is the only other parallel work on a zero-shot foundation model for time-series forecasting. However the model is currently not public access, and several model details and the benchmark dataset have not been revealed.
@@ 3 Problem Definition
The task at hand is to build a general purpose zero-shot forecaster that takes in the past $C$ time-points of a time-series as context and predicts the future $H$ time-points. Let the context be denoted by $\mathbf{y}_{1:L}:=\{y_{1},\cdots,y_{L}\}$ where we follow a numpy-like notation for indices. Similarly the actual values in the horizon is denoted by $\mathbf{y}_{L+1:L+H}$. Note that since we are building a single pre-trained model, we cannot have dataset specific dynamic or static covariates during training time. However, the datetime column is ubiquitous in all time-series data, so we can optionally have date derived features like day or the week, month of the year etc processed into a vector at each time-point $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{r}$. Such features could be available for forecasting in both the context and horizon, represented as $\mathbf{x}_{1:L+H}$. The task is then to learn a capable foundation model that can map any time-series context to horizon, given by
$f:\mathopen{}\mathclose{{}\left(\mathbf{y}_{1:L},{\bm{x}}_{1:L+H}}\right)\longrightarrow\hat{\mathbf{y}}_{L+1:L+H}.$ (1)
The accuracy of the prediction will be measured by a metric that quantifies their closeness to the actual values. For instance, if the metric is Mean Squared Error (MSE), then the goodness of fit is measured by,
$\mathrm{MSE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)$
$=\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{2}^{2}.$ (2)
@@ 4 Model Architecture
Patching. Inspired by the success of patch based modeling in the recent long horizon forecasting work [19] we also choose to breakdown the time-series into patches during training. A patch of a time-series is a natural analogues for a token in language models and have been shown to improve performance. Moreover this improves inference speed as the number of tokens being fed into the transformer is reduced by a factor of the patch length. On the other hand, increasing the patch length all the way to the context length moves us away from decoder-only training and the efficiencies that come with it. We delve into this further in Section 6.2.
@@
Decoder-only model. A key difference between our architecture and PatchTST [19] is that our model is trained in decoder-only mode [15]. In other words, given a sequence of input patches the model is optimized to predict the next patch as a function of all past patches. Similar to LLM’s this can be done in parallel over the entire context window, and automatically enables the model to predict the future after having seen varying number of input patches.
@@
Longer output patches. In LLM’s the output is always generated in an auto-regressive fashion one token at a time. However, in long-horizon forecasting it has been observed that directly predicting the full horizon yields better accuracy that multi-step auto-regressive decoding [33]. But the latter is not possible when the horizon length is not known a prior like in the case of zero-shot forecasting, which is our primary goal.
@@
We propose a middle ground by allowing the output patches for prediction to be be longer than the input patches. As an example, suppose the input patch length is 32 and output patch length is 128. During training, the model is simultaneously trained to use the first 32 time-points to forecast the next 128 time-steps, the first 64 time-points to forecast time-steps 65 to 192, the first 96 time-points to forecast time-steps 97 to 224 and so on. During inference, suppose the model is given a new time-series of length 256 and tasked with forecasting the next 256 time-steps into the future. The model will first generate the future predictions for time-steps 257 to 384, then condition on the initial 256 length input plus the generated output to generate time-steps 385 to 512. On the other hand, if in our model the output patch length was equal to the input patch length of 32, then for the same task we would have to go through 8 auto-regressive generation steps instead of just the 2 above.
@@
Patch Masking. If we use patches naively, the model might only learn to predict well for context lengths that are multiples of the input patch length. Therefore we make a careful use of masking during training. Parts of patches as well as entire patches from the beginning of the context window can be masked in a data batch. We employ a specific random masking strategy during training that helps to model see any all possible context lengths starting from 1 to a maximum context length.
@@
Now that we have mentioned the guiding principles, we next formally describe each component of our architecture that is illustrated in Figure 1.
@@
Figure 1: We provide an illustration of our model architecture during training where we show a input time-series of a certain length that can be broken down into input patches. Each patch along with (optional) time-features is processed into a vector by a residual block (as defined in the model definition) to the model dimension of the transformer layers. The vector is then added to positional encodings and fed into $n_{l}$ stacked transformer layers. SA refers to self-attention (note that we use multi-head causal attention) and FFN is the fully connected layer in the transformer. The output tokens are then mapped through a residual block to an output of size output_patch_len which is the forecast for the time-period following the last input patch seen by the model so far.
@@
Input Layers. The job of the input layers is to preprocess the time-series into input tokens to the transformer layers. We first break the input into contiguous non-overlapping patches. Then each patch is processed by a Residual Block into a vector of size model_dim. Along with the input, we also supply a binary padding mask $\mathbf{m}_{1:L}$ where $1$ denotes that the corresponding input in $\mathbf{y}_{1:L}$ should be ignored and vice-versa. The Residual Block is essentially a Multi-layer Perceptron (MLP) block with one hidden layer with a skip connection as defined in [6].
@@
Stacked Transformer. The bulk of the parameters in our model are in num_layers ($n_{l}$) transformer layers stacked on top of each other. Each of these layers have the standard multi-head self-attention (SA) followed by a feed-forward network (FFN). The main hyperparameters are model_dim which is equal to the dimension of the input tokens $\mathbf{t}_{j}$’s and number of heads (num_heads). We set the hidden size of the FFN’s to be equal to model_dim as well. We use causal attention that is each output token can only attend to input tokens that come before it in the sequence (including the corresponding input token). This can be described by the equation
$,\displaystyle\mathbf{o}_{j}=\texttt{StackedTransformer}((\mathbf{t}_{1},\dot{m}_{1}),$$\cdots,(\mathbf{t}_{j},\dot{m}_{j})),$
$\forall j\in[N].$ (4)
In Equation (4), $\dot{m}_{j}$ is the masking indicator for the $j$-th token defined as $\min\{\mathbf{m}_{p(j-1)+1:pj}\}$ i.e if a patch has any non-masked time-point the corresponding token marked as not being masked. All patches that are masked out completely are not attended to by the causal self attention.
@@
Loss Function. In this work, we focus on point forecasting. Therefore we can use a point forecasting loss during training like MSE as defined in Equation (2). The loss that is minimized during training can be expressed as,
@@
Training. We train the model with standard mini-batch gradient descent in decoder only fashion, that goes through all windows for a time-series and across time-series. The only non-standard part is the way we sample the mask during training. For each time-series in the batch, we sample a random number $r$ between 00 and $p-1$. Then we set the $\mathbf{m}_{1:r}=1$ and the rest as zero i.e we mask our a fraction of the first input patch. However, this is sufficient to cover all input context lengths from 1 to the maximum training context length.
Let us say that the maximum context length is 512 and $p=32$. Then if $r=4$, the output prediction after seeing the first patch (from $\mathbf{o}_{1}$) is optimized predict after seeing $28=32-4$ time-points, the output of the next patch (from $\mathbf{o}_{2}$) is optimized to predict after seeing $28+32$ time-points, and so on. When this argument is repeated for all such $r$’s, we can see that the model has seen all possible context lengths from $1$ to $512$.
Inference. The trained network can be used to produce forecasts for any horizon using auto-regressive decoding similar to large language models. Given an input $\mathbf{y}_{1:L}$ (assume $L$ is a multiple of $p$ for simplicity) it can first predict $\hat{\mathbf{y}}_{L+1:L+h}$. Then, we can use the concatenated vector $\tilde{\mathbf{y}}_{1:L+h}=[\mathbf{y}_{1:L};\hat{\mathbf{y}}_{L+1:L+h}]$ as an input to the network to generate the next output patch prediction $\hat{\mathbf{y}}_{L+h+1:L+2h}$ and so on. If $L$ is not a multiple of $p$ we can append some zeros to make it a multiple of $p$ and mark the corresponding entries in the mask as 1.
We name our Time-series Foundation Model as TimesFM.

### v4

@@ Abstract
Author names are listed in alphabetical order.
Motivated by recent advances in large language models for Natural Language Processing (NLP), we design a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset. Our model is based on pretraining a decoder style attention model with input patching, using a large time-series corpus comprising both real-world and synthetic datasets. Experiments on a diverse set of previously unseen forecasting datasets suggests that the model can yield accurate zero-shot forecasts across different domains, forecasting horizons and temporal granularities.
@@ 1 Introduction
Time-series data is ubiquitous in various domains such as retail, finance, manufacturing, healthcare and natural sciences. In many of these domains, one of the most important use-cases of time-series data is forecasting. Time-series forecasting is critical to several scientific and industrial applications, like retail supply chain optimization, energy and traffic prediction, and weather forecasting. In recent times, Deep learning models [30, 25] have emerged as a popular approach for forecasting rich, multivariate, time-series data, often outperforming classical statistical approaches such as ARIMA or GARCH [3]. In several forecasting competitions such as the M5 competition [22] and IARAI Traffic4cast contest [15] deep network based solutions performed very well.
@@
This motivates the question: “Can large pretrained models trained on massive amounts of time-series data learn temporal patterns that can be useful for time-series forecasting on previously unseen datasets?” In particular, can we design a time-series foundation model that obtains good zero-shot out-of-the-box forecasting performance ? Such a pretrained time-series foundation model, if possible, would bring significant benefits for downstream forecasting users in terms of no additional training burden and significantly reduced compute requirements. It is not immediately obvious that such a foundation model for time-series forecasting is possible. Unlike in NLP, there is no well defined vocabulary or grammar for time-series. Additionally, such a model would need to support forecasting with varying history lengths (context) , prediction lengths (horizon) and time granularities. Furthermore, unlike the huge volume of public text data for pretraining language models, vast amounts of time-series data is not readily available. In spite of these issues, we provide evidence to answer the above question in the affirmative.
In particular, we design TimesFM, a single foundation model for time-series forecasting that, when applied to a variety of previously-unseen forecasting datasets across different domains, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets). Our model can work well across different forecasting history lengths, prediction lengths and time granularities at inference time. The key elements of our foundation model are twofold: 1) a large-scale time-series corpus built using both real-world (mostly time-series data from web search queries [^1] https://trends.google.com and Wikipedia page visits [^2] https://wikimedia.org/api/rest_v1/) and synthetic data, which meets the volume and diversity of data needed for training our foundation model, and 2) a decoder style attention architecture with input patching, that can be efficiently pre-trained on this time-series corpus.
Compared to the latest large language models, our time-series foundation model is much smaller in both parameter size (200M parameters) and pretraining data size (O(100B) timepoints); yet we show that even at such scales, it is possible to pretrain a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time-series data. Our work also suggests that unlike recent work [11] that recommends Large Language Models such as GPT-3 and LLama-2 as out-of-the-box zero-shot forecasters, foundation models trained from scratch exclusively on time-series data can obtain much better zero-shot performance at a tiny fraction of its costs.
@@ 2 Related Work
In the last decade, deep learning models [30, 25] have emerged as powerful contenders in forecasting time-series in the presence of large training datasets and have been shown to outperform traditional statistical methods such as ARIMA and Exponential smoothing [20]. Forecasting models can be categorized broadly into: (i) Local univariate models that include traditional methods like ARIMA, exponential smoothing [20] and non-autoregressive models like Prophet [32]. These models are trained individually for each time-series in a dataset in order to predict the corresponding time-series’s future. (ii) Global univariate models like DeepAR [30], Temporal Convolutions [2], N-BEATS [25] and long-term forecasting models such as [24, 7] that are trained globally on many time-series but during inference they predict the future of a time-series as a function of its own past and other related covariates. (iii) Global multivariate models that take in the past of all time-series in the dataset to predict the future of all the time-series. Such models include the classical VAR model [43] as well as deep learning models like [31, 41, 4] to name a few.
All the works cited above have primarily been applied in the supervised setting with the notable exception of PatchTST [24] and N-BEATS [25]. PatchTST has a section on dataset-to-dataset transfer learning in the semi-supervised setting. [26] also show that the N-BEATS architecture lends itself to transfer learn between various source-target dataset pairs. However, none of these works aim to train a single foundation model that can work on a plethora of datasets. For an in-depth discussion about transfer learning in time-series we refer the reader to the survey in [21].
There has been some very recent work on re-using or fine-tuning large language models for time-series forecasting. In particular,  [11] benchmarks pretrained LLMs like GPT-3 and LLaMA-2 for zero-shot forecasting performance. As we show later, our model obtains much superior zero-shot performance at a tiny fraction of these model sizes. [42] and [6] show how to fine-tune a GPT-2 [29] backbone model for time-series forecasting tasks. With the exception of a transfer-learning study (forecasting on a target dataset after having trained on a source dataset), these papers mostly focus on fine-tuning a pretrained model on target datasets, and not on pretraining a single foundation model with good out-of-the box zero-shot performance on a variety of datasets. To the best of our knowledge, the very recent work in TimeGPT-1 [12] is the only other parallel work on a zero-shot foundation model for time-series forecasting. However the model is not public access, and several model details and the benchmark dataset have not been revealed.
@@ 3 Problem Definition
The task at hand is to build a general purpose zero-shot forecaster that takes in the past $C$ time-points of a time-series as context and predicts the future $H$ time-points. Let the context be denoted by $\mathbf{y}_{1:L}:=\{y_{1},\cdots,y_{L}\}$ where we follow a numpy-like notation for indices. Similarly the actual values in the horizon are denoted by $\mathbf{y}_{L+1:L+H}$. Note that since we are building a single pre-trained model, we cannot have dataset specific dynamic or static covariates during training time. The task is then to learn a foundation model that can map any time-series context to horizon,
$f:\mathopen{}\mathclose{{}\left(\mathbf{y}_{1:L}}\right)\longrightarrow\hat{\mathbf{y}}_{L+1:L+H}.$ (1)
The accuracy of the prediction can be measured by a metric that quantifies their closeness to the actual values, for instance, Mean Absolute Error (MAE) defined in Equation 6.
@@ 4 Model Architecture
Patching. Inspired by the success of patch based modeling in the recent long horizon forecasting work [24] we also choose to break down the time-series into patches during training. A patch of a time-series is a natural analogue for a token in language models and patching has been shown to improve performance. Moreover this improves inference speed as the number of tokens being fed into the transformer is reduced by a factor of the patch length. On the other hand, increasing the patch length all the way to the context length moves us away from decoder-only training and the efficiencies that come with it. We delve into this further in Section 6.2.
@@
Decoder-only model. A key difference between our architecture and PatchTST [24] is that our model is trained in decoder-only mode [18]. In other words, given a sequence of input patches, the model is optimized to predict the next patch as a function of all past patches. Similar to LLMs this can be done in parallel over the entire context window, and automatically enables the model to predict the future after having seen varying number of input patches.
@@
Longer output patches. In LLMs the output is always generated in an auto-regressive fashion one token at a time. However, in long-horizon forecasting it has been observed that directly predicting the full horizon yields better accuracy than multi-step auto-regressive decoding [40]. But this is not possible when the horizon length is not known apriori, as in the case of zero-shot forecasting which is our primary goal.
@@
We propose a middle ground by allowing our output patches for prediction to be longer than the input patches. As an example, suppose the input patch length is 32 and output patch length is 128. During training, the model is simultaneously trained to use the first 32 time-points to forecast the next 128 time-steps, the first 64 time-points to forecast time-steps 65 to 192, the first 96 time-points to forecast time-steps 97 to 224 and so on. During inference, suppose the model is given a new time-series of length 256 and tasked with forecasting the next 256 time-steps into the future. The model will first generate the future predictions for time-steps 257 to 384, then condition on the initial 256 length input plus the generated output to generate time-steps 385 to 512. On the other hand, if in a model the output patch length was fixed to the input patch length of 32, then for the same task we would have to go through 8 auto-regressive generation steps instead of just the 2 above. However, there is a trade-off. If the output patch length is too long, then it is difficult to handle time-series whose lengths are less than the output patch length for instance monthly, yearly time-series in our pretraining data.
@@
Patch Masking. If we use patches naively, the model might only learn to predict well for context lengths that are multiples of the input patch length. Therefore we make a careful use of masking during training. Parts of patches as well as entire patches from the beginning of the context window can be masked in a data batch. We employ a specific random masking strategy (described later) during training that helps the model see all possible context lengths starting from 1 to a maximum context length.
@@
Now that we have mentioned the guiding principles, we next formally describe each component of our model architecture (illustrated in Figure 1), which we name as TimesFM (Time-series Foundation Model).
@@
Figure 1: We provide an illustration of the TimesFM model architecture during training, where we show a input time-series of a specific length that can be broken down into input patches. Each patch along is processed into a vector by a residual block (as defined in the model definition) to the model dimension of the transformer layers. The vector is then added to positional encodings and fed into $n_{l}$ stacked transformer layers. SA refers to self-attention (note that we use multi-head causal attention) and FFN is the fully connected layer in the transformer. The output tokens are then mapped through a residual block to an output of size output_patch_len, which is the forecast for the time window following the last input patch seen by the model so far.
@@
Input Layers. The job of the input layers is to preprocess the time-series into input tokens to the transformer layers. We first break the input into contiguous non-overlapping patches. Then each patch is processed by a Residual Block into a vector of size model_dim. Along with the input, we also supply a binary padding mask $\mathbf{m}_{1:L}$ where $1$ denotes that the corresponding input in $\mathbf{y}_{1:L}$ should be ignored and vice-versa. The Residual Block is essentially a Multi-layer Perceptron (MLP) block with one hidden layer with a skip connection, similar to that defined in [7].
@@
Stacked Transformer. The bulk of the parameters in our model are in num_layers ($n_{l}$) transformer layers stacked on top of each other. Each of these layers have the standard multi-head self-attention (SA) followed by a feed-forward network (FFN). The main hyperparameters are model_dim which is equal to the dimension of the input tokens $\mathbf{t}_{j}$’s and number of heads (num_heads). We set the hidden size of the FFNs to be equal to model_dim as well. We use causal attention that is each output token can only attend to input tokens that come before it in the sequence (including the corresponding input token). This can be described by the equation
$,\displaystyle\mathbf{o}_{j}=\texttt{StackedTransformer}((\mathbf{t}_{1},\dot{m}_{1}),$$\cdots,(\mathbf{t}_{j},\dot{m}_{j})),$ (3)
for all $j\in[N]$. $\dot{m}_{j}$ is the masking indicator for the $j$-th token defined as $\min\{\mathbf{m}_{p(j-1)+1:pj}\}$ i.e if a patch has any non-masked time-point the corresponding token marked as not being masked. All patches that are masked out completely are not attended to by the causal self attention.
@@
Loss Function. In this work, we focus on point forecasting. Therefore we can use a point forecasting loss during training like Mean Squared Error (MSE). The loss that is minimized during training can be expressed as,
@@
Training. We train the model with standard mini-batch gradient descent in decoder-only fashion, that goes through all windows for a time-series and across time-series. The only non-standard part is the way we sample the mask during training. For each time-series in the batch, we sample a random number $r$ between 00 and $p-1$. Then we set the $\mathbf{m}_{1:r}=1$ and the rest as zero i.e we mask out a fraction of the first input patch. However, this is sufficient to cover all input context lengths from 1 to the maximum training context length. We explain this using an example below:
Suppose the maximum context length is 512 and $p=32$. Then if $r=4$, the output prediction after seeing the first patch (from $\mathbf{o}_{1}$) is optimized to predict after seeing $28=32-4$ time-points, the output of the next patch (from $\mathbf{o}_{2}$) is optimized to predict after seeing $28+32$ time-points, and so on. When this argument is repeated for all such $r$’s, the model has seen all possible context lengths till $512$.
Inference. The trained network can be used to produce forecasts for any horizon using auto-regressive decoding similar to large language models. Given an input $\mathbf{y}_{1:L}$ (assume $L$ is a multiple of $p$ for simplicity) it can first predict $\hat{\mathbf{y}}_{L+1:L+h}$. Then, we can use the concatenated vector $\tilde{\mathbf{y}}_{1:L+h}=[\mathbf{y}_{1:L};\hat{\mathbf{y}}_{L+1:L+h}]$ as an input to the network to generate the next output patch prediction $\hat{\mathbf{y}}_{L+h+1:L+2h}$ and so on. If $L$ is not a multiple of $p$, we simply append zeros to make it a multiple of $p$ and mark the corresponding entries in the mask as 1.